# Adapting Pretrained Language Models for Citation Classification via Self-Supervised Contrastive Learning

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extraction
@@ 2. Related Work
Additional citation analytical tasks may not necessarily originate from citation context analysis. For instance, citation prediction (Hao et al., 2024; Jin et al., 2023; Zhang et al., 2023; Cohan et al., 2020) aims to forecast potential citations from a candidate paper collection for the target document. Likewise, citation recommendation  (He et al., 2010) attempts to recommend scholarly references to the authors during the paper drafting. These analytical tasks differ more profoundly from the present research, primarily in that they focus on the stages where formal academic texts and actual citations may not yet exist.
@@ 3.2. Pretrained Language Models
As the interaction with PLMs formulated as Equation 1, our methodology only cares about how to prompt for the backbone model and to obtain the $x$ properly but is indifferent to the inner architectures. Therefore, our framework can work together with those PEFT methods (Han et al., 2024), which select, reparameterize (Yu et al., 2023), or insert (Houlsby et al., 2019) trainable modules in the PLMs to reduce trainable parameters while keeping the input and output interface unchanged, making it possible to efficiently finetune the LLMs with the limited citation data.
@@ 4.4. Complexity Analysis
Denote the number of trainable parameters in $\mathcal{M}$ as $N_{0}$, the hidden state size of $\mathcal{M}$ as $d_{x}=|x|$, the model size as $d_{z}=|z|$, the intermediate embedding size of the adapter as $d$, $C=|\mathcal{C}|$ is the size of label set, the number of trainable parameters of Citss  is $N_{0}+d(d_{x}+d_{z})+d+d_{z}+C(d_{z}+1)$. With $C\ll d_{z}$ and $d_{z}<d_{x}$, the number of parameters is in O($N_{0}+2d\cdot d_{x}$). Additionally, our method can work seamlessly with prevalent parameter-efficient finetuning methods for LLM, such as Lora, which wraps $\mathcal{M}$ and reduces $N_{0}$ without changing its interface with our framework. As for time complexity, since SC only shortens the $T_{i}$ and KP performs replacement between synonyms or STKs, the length of $\tilde{T_{i}}$ after tokenization resembles that of $T_{i}$, implying similar forwarding complexity. Assume the PLM forwarding complexity per sample is $O(|\mathcal{M}|)$, the amortized complexity to perform contrastive learning amongst the batch is $O(|\mathcal{B}|\cdot d_{z}^{2})$, hence the training complexity of our framework per sample is $O(3|\mathcal{M}|+6d\cdot d_{x}+2|\mathcal{B}|\cdot d_{z}^{2})\approx O(3|\mathcal{M}|)$ and the inference complexity is $O(|\mathcal{M}|+2d\cdot d_{x})\approx O(|\mathcal{M}|)$. Here the time for for SC and KP can be neglected, because with $l^{\prime}$ be the length of $T_{i}$, it takes $O(l^{\prime})$ time to scan the context and generate a transformed sample, while $|\mathcal{M}|$ is generally in O($l^{\prime 2}$).
@@ 5.1.3. Baselines
We then introduce 3 baselines that are dedicated to finetune encoder-based PLM. (2) TRL (Shui et al., 2024): It is a multi-task learning framework that uses the labeled data from auxiliary datasets to aid the finetuning on the primary dataset. A task relation learning procedure automatically computes the task weights. In our experiments, we use ACL-ARC and FOCAL as the auxiliary datasets for each other; for ACT2, neither of the other datasets is helpful so we only use itself. (3) IREL (Maheshwari et al., 2021): It is the winning system in the 2021 SDP Citation Context Classification Shared task (Kunnath et al., 2021b), which finetune SciBERT and a linear classifier end-to-end with a class-balanced classification loss (King and Zeng, 2001). (4) PET (Kunnath et al., 2023): It explored several closed-form prompts to finetune SciBERT in the Pattern Explicit Tuning style (Ahmad et al., 2017). It reports P1 as one of the best prompts across different datasets, resulting in previously state-of-the-art performance.
@@ 5.2.1. Classification Performance
The overall comparison is reported in Table 2. Macro-f1 represents the average performance on each class since the citation types are unevenly distributed as in Table 1. Based on the results, we have the following four main observations. (1) Citss  achieves the state-of-the-art performance, outperforming existing methods on 5 out of 6 metrics when using SciBERT and on all metrics when using the Llama3-8B backbone. This highlights the versatility and effectiveness of our approach with both encoder-based PLMs and decoder-based LLMs. On ACT2, it is hard to achieve high scores on both metrics due to the difficulties in predicting its minority classes. IREL prioritizes Macro-F1 by weighting minority classes more heavily, compared to which our framework strikes a better balance and achieves significantly higher accuracy than IREL without substantial loss in Macro-F1. (2) Turning to the experiments with SciBERT, PET achieves the second-best overall performance. The success of both our method and PET, which leverage prompting strategies, highlights the important role of task-specific patterns in adapting PLMs for citation classification. TRL attains the second-best accuracy on FOCAL and ACT2, albeit for different reasons. On FOCAL, its performance validates the effectiveness of supplementing labeled data from the auxiliary ACL-ARC dataset. However, it becomes a dummy classifier on ACT2 that overwhelmingly predicts the majority class, resulting in the worst Macro-F1 score. (3) In experiments utilizing decoder-based LLMs, the finetuning methods essentially outperform the IFP. This superiority extends even when comparing finetuning to IFP using the significantly larger Instruction-tuned 70B model. This reflects that there is still a large gap between the generation approach, which relies solely on task descriptions and examples, and the data-driven finetuning approaches of cutting-edge LLMs for the complex text classification task of citation classification. (4) With Citss, Llama3-8B outperforms SciBERT on ACL-ARC, achieving significantly better results. While the performance gap between Llama3-8B and SciBERT narrows on Focal and ACT2, this suggests that Focal and ACT2 may benefit more from scientific text-specialized pretraining of SciBERT. ACL-ARC appears to leverage Llama3-8B’s larger scale pretraining on general domain knowledge.
@@ 5.2.2. Efficiency
We now detail the absolute time costs under our experiment environments. (1) For Citss, with SciBERT, it takes 17 to 23 minutes for training and 3 to 12 seconds for inference on the entire test set. With Llama3-8B and LoRA, it takes around 89, 160, and 131 minutes for training and 12, 71, and 53 seconds for inference on ACL-ARC, FOCAL, and ACT datasets, respectively. The time for producing transformed samples with KP and SC is negligible in comparison. (2) Among the baselines, the feature-based Scaffold is the most efficient, requiring only a few minutes for training and seconds for inference. However, this efficiency comes at the cost of significant manual effort in data preprocessing, particularly in collecting and cleaning the section name feature. (3) All SciBERT-based baselines share similar inference times. As for the training, IREL and PET require similar training time, which is around 6 to 12 minutes. TRL involves augmenting the training data and determining auxiliary task weights, resulting in a comparable overall training time to Citss with SciBERT. (4) For LLM baselines, while IFP is training-free, its inference speed is significantly slower due to the sequential decoding, requiring roughly 5 and 15 seconds per sample for Llama3-8B and Llama3-70B, respectively. This limits its practical applicability. LoRA finetuning, on the other hand, takes 40 to 92 minutes for training and 11 to 52 seconds for inference.
@@ 6. Conclusion
We introduce a framework, Citss, to finetune PLMs for citation classification via self-supervised contrastive learning. Our framework employs sentence-level cropping and keyphrase perturbation strategies to construct contrastive pairs without the need for labels. Taking the citation context as input, our framework acquires task-specific representation from the output of PLMs, and performs contrastive learning together with the supervised loss over the labeled data. Experiments with three benchmark datasets demonstrated that our framework improves the classification performance by an average 3.9% with SciBERT and 6.3% with Llama3-8B, compared to the previous state of the art.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736829isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Information extraction
@@ 2. Related Work
Additional citation analytical tasks may not necessarily originate from citation context analysis. For instance, citation prediction (Hao et al., 2024; Jin et al., 2023; Zhang et al., 2023; Cohan et al., 2020) aims to forecast potential citations from a candidate paper collection for the target document. Likewise, citation recommendation  (He et al., 2010) attempts to recommend scholarly references to the authors during the paper drafting. These tasks differ more profoundly from the present research, primarily in that they focus on the stages where formal academic texts and actual citations may not yet exist.
@@ 3.2. Pretrained Language Models
As the interaction with PLMs formulated as Equation 1, our methodology only cares about how to prompt for the backbone model and to obtain the $x$ properly but is indifferent to the inner architectures. Therefore, our framework can work together with those PEFT methods (Han et al., 2024), which select, reparameterize (Yu et al., 2023), or insert (Houlsby et al., 2019) trainable modules in the PLMs to reduce trainable parameters while keeping the input and output interface unchanged, making it possible to efficiently fine-tune the LLMs with the limited citation data.
@@ 4.4. Complexity Analysis
Denote the number of trainable parameters in $\mathcal{M}$ as $N_{0}$, the hidden state size of $\mathcal{M}$ as $d_{x}=|x|$, the model size as $d_{z}=|z|$, the intermediate embedding size of the adapter as $d$, $C=|\mathcal{C}|$ is the size of label set, the number of trainable parameters of Citss  is $N_{0}+d(d_{x}+d_{z})+d+d_{z}+C(d_{z}+1)$. With $C\ll d_{z}$ and $d_{z}<d_{x}$, the number of parameters is in O($N_{0}+2d\cdot d_{x}$). Additionally, our method can work seamlessly with prevalent parameter-efficient fine-tuning methods for LLM, such as Lora, which wraps $\mathcal{M}$ and reduces $N_{0}$ without changing its interface with our framework. As for time complexity, since SC only shortens the $T_{i}$ and KP performs replacement between synonyms or STKs, the length of $\tilde{T_{i}}$ after tokenization resembles that of $T_{i}$, implying similar forwarding complexity. Assume the PLM forwarding complexity per sample is $O(|\mathcal{M}|)$, the amortized complexity to perform contrastive learning amongst the batch is $O(|\mathcal{B}|\cdot d_{z}^{2})$, hence the training complexity of our framework per sample is $O(3|\mathcal{M}|+6d\cdot d_{x}+2|\mathcal{B}|\cdot d_{z}^{2})\approx O(3|\mathcal{M}|)$ and the inference complexity is $O(|\mathcal{M}|+2d\cdot d_{x})\approx O(|\mathcal{M}|)$. Here the time for for SC and KP can be neglected, because with $l^{\prime}$ be the length of $T_{i}$, it takes $O(l^{\prime})$ time to scan the context and generate a transformed sample, while $|\mathcal{M}|$ is generally in O($l^{\prime 2}$).
@@ 5.1.3. Baselines
We then introduce 3 baselines that are dedicated to finetune encoder-based PLM. (2) TRL (Shui et al., 2024): It is a multi-task learning framework that uses the labeled data from auxiliary datasets to aid the fine-tuning on the primary dataset. A task relation learning procedure automatically computes the task weights. In our experiments, we use ACL-ARC and FOCAL as the auxiliary datasets for each other; for ACT2, neither of the other datasets is helpful so we only use itself. (3) IREL (Maheshwari et al., 2021): It is the winning system in the 2021 SDP Citation Context Classification Shared task (Kunnath et al., 2021b), which fine-tunes SciBERT and a linear classifier end-to-end with a class-balanced classification loss (King and Zeng, 2001). (4) PET (Kunnath et al., 2023): It explored several closed-form prompts to fine-tune SciBERT in the Pattern Explicit Tuning style (Ahmad et al., 2017). It reports P1 as one of the best prompts across different datasets, resulting in previously state-of-the-art performance.
@@ 5.2.1. Classification Performance
The overall comparison is reported in Table 2. Macro-f1 represents the average performance on each class since the citation types are unevenly distributed as in Table 1. Based on the results, we have the following four main observations. (1) Citss  achieves the state-of-the-art performance, outperforming existing methods on 5 out of 6 metrics when using SciBERT and on all metrics when using the Llama3-8B backbone. This highlights the versatility and effectiveness of our approach with both encoder-based PLMs and decoder-based LLMs. On ACT2, it is hard to achieve high scores on both metrics due to the difficulties in predicting its minority classes. IREL prioritizes Macro-F1 by weighting minority classes more heavily, compared to which our framework strikes a better balance and achieves significantly higher accuracy than IREL without substantial loss in Macro-F1. (2) Turning to the experiments with SciBERT, PET achieves the second-best overall performance. The success of both our method and PET, which leverage prompting strategies, highlights the important role of task-specific patterns in adapting PLMs for citation classification. TRL attains the second-best accuracy on FOCAL and ACT2, albeit for different reasons. On FOCAL, its performance validates the effectiveness of supplementing labeled data from the auxiliary ACL-ARC dataset. However, it becomes a dummy classifier on ACT2 that overwhelmingly predicts the majority class, resulting in the worst Macro-F1 score. (3) In experiments utilizing decoder-based LLMs, the fine-tuning methods essentially outperform the IFP. This superiority extends even when comparing fine-tuning to IFP using the significantly larger Instruction-tuned 70B model. This reflects that there is still a large gap between the generation approach, which relies solely on task descriptions and examples, and the data-driven fine-tuning approaches of cutting-edge LLMs for the complex text classification task of citation classification. (4) With Citss, Llama3-8B outperforms SciBERT on ACL-ARC, achieving significantly better results. While the performance gap between Llama3-8B and SciBERT narrows on Focal and ACT2, this suggests that Focal and ACT2 may benefit more from scientific text-specialized pretraining of SciBERT. ACL-ARC appears to leverage Llama3-8B’s larger scale pretraining on general domain knowledge.
@@ 5.2.2. Efficiency
We now detail the absolute time costs under our experiment environments. (1) For Citss, with SciBERT, it takes 17 to 23 minutes for training and 3 to 12 seconds for inference on the entire test set. With Llama3-8B and LoRA, it takes around 89, 160, and 131 minutes for training and 12, 71, and 53 seconds for inference on ACL-ARC, FOCAL, and ACT datasets, respectively. The time for producing transformed samples with KP and SC is negligible in comparison. (2) Among the baselines, the feature-based Scaffold is the most efficient, requiring only a few minutes for training and seconds for inference. However, this efficiency comes at the cost of significant manual effort in data preprocessing, particularly in collecting and cleaning the section name feature. (3) All SciBERT-based baselines share similar inference times. As for the training, IREL and PET require similar training time, which is around 6 to 12 minutes. TRL involves augmenting the training data and determining auxiliary task weights, resulting in a comparable overall training time to Citss with SciBERT. (4) For LLM baselines, while IFP is training-free, its inference speed is significantly slower due to the sequential decoding, requiring roughly 5 and 15 seconds per sample for Llama3-8B and Llama3-70B, respectively. This limits its practical applicability. LoRA fine-tuning, on the other hand, takes 40 to 92 minutes for training and 11 to 52 seconds for inference.
@@ 6. Conclusion
We introduce a framework, Citss, to fine-tune PLMs for citation classification via self-supervised contrastive learning. Our framework employs sentence-level cropping and keyphrase perturbation strategies to construct contrastive pairs without the need for labels. Taking the citation context as input, our framework acquires task-specific representation from the output of PLMs, and performs contrastive learning together with the supervised loss over the labeled data. Experiments with three benchmark datasets demonstrated the superiority of our framework.
Acknowledgements.
Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab and 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA. Yongqi Zhang’s work is supported by Guangdong Basic and Applied Basic Research Foundation 2025A1515010304, and Guangzhou Science and Technology Planning Project 2025A03J4491.